# Topological symmetry groups of the Generalized Petersen Graphs

### v1

@@ 7. The exceptional case $P(10,2)$
Frucht proved that $\mathrm{Aut}(P(10,2))=A_{5}\times\mathbb{Z}_{2}$ [12]. The embedding $\Gamma$ of $P(10,2)$ as the 1-skeleton of a regular dodecahedron (see the left image of 7) will be the basis for all of the embeddings in this section. The group $G=A_{5}$ of rotations of a solid dodecahedron consists of six order $5$ rotations about an axis through the centers of opposite faces, ten order $3$ rotations about an axis through opposite vertices, and 15 order $2$ rotations about an axis through midpoints of opposite edges.
Figure 7. On the left, $\Gamma$ is the $1$-skeleton of a dodecahedron. On the right are red and blue inscribed tetrahedra which will be used in the proof of Proposition 7.3.
The automorphism $\alpha=(v_{2}v_{8})(u_{2}u_{8})(v_{6}v_{4})(u_{6}u_{4})(v_{3}v_{7})(u_{3}u_{7})(v_{1}v_{9})(u_{1}u_{9})$ is not induced by a rotation of the solid dodecahedron, and hence is not in $G$. However, it is induced by a rotation $h$ of $S^{3}$ around the equator of $\Gamma$ containing $\overline{u_{10}v_{10}}$ and $\overline{u_{5}v_{5}}$ which interchanges the inside and outside of the dodecahedron. Thus $H=\langle G,h\rangle=A_{5}\times\mathbb{Z}_{2}$ is a group of rotations of $(S^{3},\Gamma)$ which induces $\mathrm{Aut}(P(10,2))=A_{5}\times\mathbb{Z}_{2}$. Hence we have proven the following.
@@
Now we modify the embedding $\Gamma$ by adding the $4_{1}$ knot to each edge which has a vertex in $T_{1}\cup T_{2}$ to get an embedding $\Gamma_{A42}$ which is invariant under $H^{\prime}$. Thus $A_{4}\times\mathbb{Z}_{2}\leq\mathrm{TSG}_{+}(\Gamma_{A42})$, and every homeomorphism of $(S^{3},\Gamma_{A42})$ leaves $T_{1}\cup T_{2}$ setwise invariant. Now observe that no automorphism of order $5$ of $P(10,2)$ leaves $T_{1}\cup T_{2}$ setwise invariant. Thus no order $5$ automorphism is in $\mathrm{TSG}_{+}(\Gamma_{A42})$. Since $A_{4}\times\mathbb{Z}_{2}$ is a maximal subgroup of $A_{5}\times\mathbb{Z}_{2}$, it follows that $A_{4}\times\mathbb{Z}_{2}=\mathrm{TSG}_{+}(\Gamma_{A42})$ induced by $H^{\prime}$.
@@
We again start with the embedding $\Gamma$ of $P(10,2)$. Now we add $4_{1}$ knots to the six edges containing $u_{1}$ or $u_{6}$ to obtain an embedding $\Gamma_{D6}$. Let $h^{\prime}$ be the rotation of $(S^{3},\Gamma_{D6})$ obtained as the composition of the rotation $h$ of order $2$ around the equator of $\Gamma$ containing $\overline{u_{10}v_{10}}$ and $\overline{u_{5}v_{5}}$ and a rotation of order $2$ around an axis that passes through the midpoints of $\overline{u_{1}u_{2}}$ and $\overline{u_{6}u_{7}}$. Then $h^{\prime}$ induces the antipodal automorphism
$\alpha^{\prime}=(u_{1}u_{6})(u_{2}u_{7})(u_{3}u_{8})(u_{4}u_{9})(u_{5}u_{10})(v_{1}v_{6})(v_{2}v_{7})(v_{3}v_{8})(v_{4}v_{9})(v_{5}v_{10}).$
Let $f$ be an order $3$ rotation of $(S^{3},\Gamma_{D6})$ around an axis that passes through the vertices $u_{1}$ and $u_{6}$. Then $f$ induces the automorphism
$\beta=(u_{2},u_{10},v_{1})(u_{3},v_{10},v_{9})(u_{4},v_{8},v_{7})(u_{5},v_{6},u_{7})(u_{8},v_{5},v_{4})(u_{9},v_{3},v_{2}).$
Finally, let $g$ be the order $2$ rotation around an axis that passes through the midpoints of edges $\overline{u_{3}u_{4}}$ and $\overline{u_{8}u_{9}}$. Then $g$ induces the automorphism
$\gamma=(u_{1},u_{6})(u_{2},u_{5})(u_{3},u_{4})(u_{7},u_{10})(u_{8},u_{9})(v_{1},v_{6})(v_{2},v_{5})(v_{3},v_{4})(v_{7},v_{10})(v_{8},v_{9}).$
Now $\langle h,g,f\rangle$ induces the group $\langle\alpha^{\prime},\beta,\gamma\rangle$ on $\Gamma_{D6}$. Since no non-trivial orientation preserving isometry of $S^{3}$ can pointwise fix $\Gamma_{D6}$ or a triad of edges sharing a vertex, the group $\langle h,g,f\rangle$ is isomorphic to $\langle\alpha^{\prime},\beta,\gamma\rangle$ and also induces an isomorphic group on the triads of edges containing $u_{1}$ or $u_{6}$. However, the action of $\langle\alpha^{\prime},\beta,\gamma\rangle$ on the two triads is $D_{6}$ because $\alpha^{\prime}$ interchanges the two triads, $\beta$ rotates each triad, and $\alpha^{\prime}\gamma$ flips each triad over fixing one edge of each. Thus $D_{6}\leq\mathrm{TSG}_{+}(\Gamma_{D6})$. On the other hand, because of the $4_{1}$ knots, not every element of $\operatorname{Aut}(P(10,2))$ can be induced on the embedding $\Gamma_{D6}$. It follows that $\mathrm{TSG}_{+}(\Gamma_{D6})$ is a proper subgroup of $\operatorname{Aut}(P(10,2))$. Since $D_{6}$ a maximal in $A_{5}\times\mathbb{Z}_{2}$, this means that $D_{6}=\mathrm{TSG}_{+}(\Gamma_{D6})$. ∎
@@ 8. The exceptional case $P(10,3)$
$\alpha$ $=(u_{1},u_{2},u_{3},u_{4},u_{5},u_{6},u_{7},u_{8},u_{9},u_{10})(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10})$
$\beta$ $=(u_{1},v_{4})(u_{2},u_{4})(u_{5},v_{2})(u_{6},v_{9})(u_{7},u_{9})(u_{10},v_{7}).$
We will see that all of the subgroups of $\textrm{Aut}(P(10,3))$ are realizable, but not all are positively realizable. We begin by determining which subgroups of $\textrm{Aut}(P(10,3))$ are not positively realizable.
@@
All order $6$ automorphisms that contain $3$-cycles are conjugate to the automorphism $\rho=(u_{1},v_{2},u_{3})(u_{4},u_{10},v_{5},v_{3},v_{1},v_{9})(u_{5},v_{10},v_{8},v_{6},v_{4},u_{9})(u_{6},v_{7},u_{8})$. If $\rho$ were positively realizable, then $\rho^{3}$ would be as well. But $\rho^{3}$ is an order $2$ automorphism that contains only six $2$-cycles, which we saw is not positively realizable.∎
@@
Below we determine the positively realizable subgroups of $\mathrm{Aut}(P(10,3))$. By Theorem 2.1 we have the following.
@@
$\nu_{1}$ $=(u_{1},u_{5})(u_{2},v_{5},v_{1},u_{4})(u_{3},v_{2},v_{8},v_{4})(u_{6},u_{10})(u_{7},v_{10},v_{6},u_{9})(u_{8},v_{7},v_{3},v_{9})$
$\nu_{2}$ $=(u_{1},u_{6})(u_{2},u_{7})(u_{3},v_{7})(u_{4},v_{10})(u_{5},u_{10})(u_{8},v_{2})(u_{9},v_{5})(v_{1},v_{6})(v_{3},v_{4})(v_{8},v_{9}).$
Let $\Gamma_{D4}$ be the embedding of $P(10,3)$ in an unknotted isometric solid torus $T$ in $S^{3}$, illustrated in Figure 8, where the edges which are on $\partial T$ are green on the flat torus and the edges of $\Gamma_{D4}$ which go into the interior of $T$ are red. The edges $\overline{u_{10}u_{1}}$ and $\overline{u_{5}u_{6}}$ are diametrically opposed arcs on the core of the solid torus illustrated in the center. The vertices in a given orbit of $\nu_{1}$ are the same color. Two special meridians are colored purple and orange on the right.
Figure 8. On the left is $\Gamma_{D4}$ on the flat torus $\partial T$, in the center is the core of the solid torus $T$, on the right are two special meridians.
@@
Figure 9. The embedding $\Gamma_{D6}$, where $v_{10}$ is at the point at $\infty$.
@@
Proposition 8.7.
Proof.
Figure 10. The $4$-symplex on the left is $\Omega$, the red graph in the center is the embedding $\Lambda$ of $P(10,3)$ in an unfolded picture of $\Omega$, on the right is the dual $4$-simplex $\Omega^{\prime}$.
Since $h$ commutes with every element of the group $G$ of isometries, $H=\langle G,h\rangle=A_{5}\times\mathbb{Z}_{2}$. Also, by our construction, $H$ leaves $\Lambda$ setwise invariant. Hence $A_{5}\times\mathbb{Z}_{2}\leq\mathrm{TSG}_{+}(\Lambda)$. Now since $A_{5}\times\mathbb{Z}_{2}$ is a maximal subgroup of $S_{5}\times\mathbb{Z}_{2}$ and $S_{5}\times\mathbb{Z}_{2}$ is not positively realizable for $P(10,3)$, we have $A_{5}\times\mathbb{Z}_{2}=\mathrm{TSG}_{+}(\Lambda)$. Thus $A_{5}\times\mathbb{Z}_{2}$ is positively realizable.
@@
Recall from the proof of Proposition 8.8 that $\Lambda_{T}$ was obtained from $\Lambda$ by adding a $4_{1}$ knot to the edges going between the tetrahedron $\Omega_{T}$ with vertices $1,2,3,4$ and the dual tetrahedron $\Omega^{\prime}_{T}$ with vertices $1245,1345,1235,1245$. We saw there that $\mathrm{TSG}_{+}(\Lambda_{T})$ is induced by the group $A_{4}\times\mathbb{Z}_{2}$ of orientation preserving isometries of $(S^{3},\Omega_{T}\cup\Omega^{\prime}_{T})$. Now observe that there is a reflection of $(S^{3},\Omega_{T}\cup\Omega^{\prime}_{T})$ which pointwise fixes the sphere containing the vertices $u_{9},v_{10},u_{1},u_{5}$ and $u_{8},v_{7},u_{2},u_{6}$. Thus the full group of isometries of $(S^{3},\Omega_{T}\cup\Omega^{\prime}_{T})$ including reflections is $S_{4}\times\mathbb{Z}_{2}$. Since these isometries induce a faithful action on $\Lambda_{T}$, it follows that $S_{4}\times\mathbb{Z}_{2}\leq\mathrm{TSG}(\Lambda_{T})$.
Let $\alpha=(u_{1},u_{2},u_{3},u_{4},u_{5},u_{6},u_{7},u_{8},u_{9},u_{10})(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},v_{9},v_{10})$. Then $\alpha\in\mathrm{Aut}(P(10,3))$ taking the edge $\overline{v_{1}v_{8}}$ to the edge $\overline{v_{2}v_{9}}$. However, in the embedding $\Lambda_{T}$, the edge $\overline{v_{1}v_{8}}$ contains a knot while the edge $\overline{v_{2}v_{9}}$ does not. Thus $\alpha$ is not contained in $\mathrm{TSG}(\Lambda_{T})$, and hence $\mathrm{TSG}(\Lambda_{T})\not=\mathrm{Aut}(P(10,3))=S_{5}\times\mathbb{Z}_{2}$. Since, $S_{4}\times\mathbb{Z}_{2}$ is a maximal subgroup of $S_{5}\times\mathbb{Z}_{2}$ it follows that $\mathrm{TSG}(\Lambda_{T})=S_{4}\times\mathbb{Z}_{2}$.
@@
We again start with the embedding $\Lambda_{T}$, but now we replace the $4_{1}$ knots on the edges $\overline{u_{7}v_{7}}$ and $\overline{u_{2}v_{2}}$ by the achiral and invertible knot $5_{2}$ to obtain an embedding $\Gamma_{D4}$ such that $\mathrm{TSG}(\Gamma_{D4})$ leaves $\{\overline{v_{7}u_{7}},\overline{u_{2}v_{2}}\}$ and $\{\overline{v_{8}v_{1}},\overline{u_{9}u_{10}},\overline{v_{6}v_{3}},\overline{u_{5}u_{4}}\}$ setwisse invariant. Then any homeomorphism of $(S^{3},\Omega_{T}\cup\Omega_{T}^{\prime})$ which leaves $\Gamma_{D4}$ setwise invariant either interchanges the squares $S=\overline{1234}\subseteq\Omega_{T}$ and
@@
$\gamma_{1}$ $=(u_{5},v_{8},u_{9},v_{6})(u_{1},v_{10},u_{3},v_{4})(u_{7},v_{2})(v_{5},u_{8},v_{9},u_{6})(v_{1},u_{10},v_{3},u_{4})(u_{2},v_{7})$
$\gamma_{2}$ $=(u_{1},v_{10})(u_{2},v_{7})(u_{3},v_{4})(u_{6},v_{5})(u_{7},v_{2})(u_{8},v_{9})(v_{1},v_{3})(v_{6},v_{8})$
$\gamma_{3}$ $=(v_{3},v_{8})(v_{1},v_{6})(v_{9},v_{4})(v_{7},v_{2})(v_{5},v_{10})(u_{3},u_{8})(u_{1},u_{6})(u_{9},u_{4})(u_{7},u_{2})(u_{5},u_{10})$

### v2

@@ 7. The exceptional case $P(10,2)$
Frucht proved that $\mathrm{Aut}(P(10,2))=A_{5}\times\mathbb{Z}_{2}$ [12]. The embedding $\Gamma$ of $P(10,2)$ as the 1-skeleton of a regular dodecahedron (see the left image of 8) will be the basis for all of the embeddings in this section. The group $G=A_{5}$ of rotations of a solid dodecahedron consists of six order $5$ rotations about an axis through the centers of opposite faces, ten order $3$ rotations about an axis through opposite vertices, and 15 order $2$ rotations about an axis through midpoints of opposite edges.
Figure 8. On the left, $\Gamma$ is an embedding of $P(10,2)$ with $\mathrm{TSG}_{+}(\Gamma)=A_{5}\times\mathbb{Z}_{2}$. On the right, red and blue inscribed tetrahedra which will be used in the proof of Proposition 7.3.
The automorphism $\alpha=(v_{2}v_{8})(u_{2}u_{8})(v_{6}v_{4})(u_{6}u_{4})(v_{3}v_{7})(u_{3}u_{7})(v_{1}v_{9})(u_{1}u_{9})$ is not induced by a rotation of the solid dodecahedron, and hence is not in $G$. However, it is induced by a rotation $h$ of $S^{3}$ around the equator of $\Gamma$ containing $\overline{u_{10}v_{10}}$ and $\overline{u_{5}v_{5}}$ which interchanges the inside and outside of the dodecahedron. Thus $H=\langle G,h\rangle=A_{5}\times\mathbb{Z}_{2}$ is a group of rotations of $(S^{3},\Gamma)$ which induces $\mathrm{Aut}(P(10,2))=A_{5}\times\mathbb{Z}_{2}$. Hence we have proved the following.
@@
Now we modify the embedding $\Gamma$ by adding the $4_{1}$ knot to each edge which has a vertex in $T_{1}\cup T_{2}$ to get an embedding $\Gamma_{A42}$ which is invariant under $H^{\prime}$. Thus $A_{4}\times\mathbb{Z}_{2}\leq\mathrm{TSG}_{+}(\Gamma_{A42})$, and every homeomorphism of $(S^{3},\Gamma_{A42})$ leaves $T_{1}\cup T_{2}$ setwise invariant. Now observe that no automorphism of order $5$ of $P(10,2)$ leaves $T_{1}\cup T_{2}$ setwise invariant. Thus no order $5$ automorphism is in $\mathrm{TSG}_{+}(\Gamma_{A42})$. Since $A_{4}\times\mathbb{Z}_{2}$ is a maximal subgroup of $A_{5}\times\mathbb{Z}_{2}$, we have $A_{4}\times\mathbb{Z}_{2}=\mathrm{TSG}_{+}(\Gamma_{A42})$ induced by $H^{\prime}$.
@@
We again start with the embedding $\Gamma$ of $P(10,2)$ from Figure 8. Let $h_{1}$ be a rotation of $(S^{3},\Gamma)$ of order $2$ around the equator of $\Gamma$ containing the edges $\overline{u_{10}v_{10}}$ and $\overline{u_{5}v_{5}}$. Then $h_{1}$ induces the automorphism
$\alpha_{1}=(u_{1}u_{9})(u_{2}u_{8})(u_{3}u_{7})(u_{4}u_{6})(v_{1}v_{9})(v_{2}v_{8})(v_{3}v_{7})(v_{4}v_{6}).$
Let $h_{2}$ be a rotation of $(S^{3},\Gamma)$ of order $2$ around an axis that passes through the midpoints of the edges $\overline{u_{7}u_{8}}$ and $\overline{u_{2}u_{3}}$. Then $h_{2}$ induces the automorphism
$\alpha_{2}=(u_{2}u_{3})(u_{6}u_{9})(u_{7}u_{8})(u_{5}u_{10})(u_{4}u_{1})(v_{2}v_{3})(v_{6}v_{9})(v_{7}v_{8})(v_{5}v_{10})(v_{4}v_{1}).$
It follows that $h=h_{2}h_{1}$ is an isometry of $(S^{3},\Gamma)$ inducing the automorphism
$\alpha^{\prime}=\alpha_{2}\alpha_{1}=(u_{1}u_{6})(u_{2}u_{7})(u_{3}u_{8})(u_{4}u_{9})(u_{5}u_{10})(v_{1}v_{6})(v_{2}v_{7})(v_{3}v_{8})(v_{4}v_{9})(v_{5}v_{10}).$
Let $f$ be an order $3$ rotation of $(S^{3},\Gamma)$ around an axis that passes through vertices $u_{1}$ and $u_{6}$. Then $f$ induces the automorphism
$\beta=(u_{2}u_{10}v_{1})(u_{3}v_{10}v_{9})(u_{4}v_{8}v_{7})(u_{5}v_{6}u_{7})(u_{8}v_{5}v_{4})(u_{9}v_{3}v_{2}).$
Finally, let $g$ be the order $2$ rotation of $(S^{3},\Gamma)$ around an axis that passes through the midpoints of edges $\overline{u_{3}u_{4}}$ and $\overline{u_{8}u_{9}}$. Then $g$ induces the automorphism
$\gamma=(u_{1}u_{6})(u_{2}u_{5})(u_{3}u_{4})(u_{7}u_{10})(u_{8}u_{9})(v_{1}v_{6})(v_{2}v_{5})(v_{3}v_{4})(v_{7}v_{10})(v_{8}v_{9}).$
Note that the pair $\{u_{1},u_{6}\}$ is setwise fixed by $\alpha^{\prime}$, $\beta$, and $\gamma$. Now we add $4_{1}$ knots to the six edges containing $u_{1}$ or $u_{6}$ to obtain an embedding $\Gamma_{D6}$ such that the isometries $h$, $g$, and $f$ leave $\Gamma_{D6}$ setwise invariant. Then $\langle h,g,f\rangle$ induces $\langle\alpha^{\prime},\beta,\gamma\rangle$ on $\Gamma_{D6}$. Since no non-trivial finite order orientation preserving isometry of $S^{3}$ can pointwise fix $\Gamma_{D6}$, the isometry group $\langle h,g,f\rangle$ is isomorphic to the automorphism group $\langle\alpha^{\prime},\beta,\gamma\rangle$. Furthermore, because of the $4_{1}$ knots on the edges of the triads centered at $u_{1}$ and $u_{6}$, every element of $\langle h,g,f\rangle$ leaves this pair of triads setwise invariant. Since no non-trivial finite order orientation preserving isometry can pointwise fix a triad, $\langle h,g,f\rangle$ induces an isomorphic action on this pair of triads.
However, the action of $\langle\alpha^{\prime},\beta,\gamma\rangle$ on the two triads is $D_{6}$ because $\alpha^{\prime}$ interchanges the two triads, $\beta$ rotates both triads, and $\alpha^{\prime}\gamma$ flips over each triad fixing the edges $\overline{u_{6}v_{6}}$ and $\overline{u_{1}v_{1}}$. Thus $D_{6}\leq\mathrm{TSG}_{+}(\Gamma_{D6})$. On the other hand, because of the $4_{1}$ knots, not every element of $\operatorname{Aut}(P(10,2))$ can be induced on the embedding $\Gamma_{D6}$. It follows that $\mathrm{TSG}_{+}(\Gamma_{D6})$ is a proper subgroup of $\operatorname{Aut}(P(10,2))=A_{5}\times\mathbb{Z}_{2}$. Since $D_{6}$ a maximal in $A_{5}\times\mathbb{Z}_{2}$, this means that $D_{6}=\mathrm{TSG}_{+}(\Gamma_{D6})$. ∎
@@ 8. The exceptional case $P(10,3)$
$\alpha$ $=(u_{1}u_{2}u_{3}u_{4}u_{5}u_{6}u_{7}u_{8}u_{9}u_{10})(v_{1}v_{2}v_{3}v_{4}v_{5}v_{6}v_{7}v_{8}v_{9}v_{10})$
$\beta$ $=(u_{1}v_{4})(u_{2}u_{4})(u_{5}v_{2})(u_{6}v_{9})(u_{7}u_{9})(u_{10}v_{7}).$
The isomorphism classes of proper nontrivial subgroups of $S_{5}\times\mathbb{Z}_{2}$ are $A_{5}\times\mathbb{Z}_{2},S_{5},A_{5},S_{4}\times\mathbb{Z}_{2},(\mathbb{Z}_{5}\rtimes\mathbb{Z}_{2})\times\mathbb{Z}_{2},S_{3}\times\mathbb{Z}_{2}^{2},A_{4}\times\mathbb{Z}_{2},S_{4},D_{10},\mathbb{Z}_{5}\rtimes\mathbb{Z}_{2},D_{4}\times\mathbb{Z}_{2},\mathbb{Z}_{6}\times\mathbb{Z}_{2},D_{6},A_{4},\mathbb{Z}_{10},D_{5},\mathbb{Z}_{2}^{3},D_{4},\mathbb{Z}_{4}\times\mathbb{Z}_{2},\mathbb{Z}_{6},S_{3},\mathbb{Z}_{5},\mathbb{Z}_{2}^{2},\mathbb{Z}_{4},\mathbb{Z}_{3}$ and $\mathbb{Z}_{2}$. We will see that all of these subgroups are realizable, but not all are positively realizable for $P(10,3)$. We begin by determining which subgroups are not positively realizable.
@@
All order $6$ automorphisms that contain $3$-cycles are conjugate to the automorphism $\rho=(u_{1},v_{2},u_{3})(u_{4}u_{10}v_{5}v_{3}v_{1}v_{9})(u_{5}v_{10}v_{8}v_{6}v_{4}u_{9})(u_{6}v_{7}u_{8})$. If $\rho$ were positively realizable, then $\rho^{3}$ would be as well. But $\rho^{3}$ is an order $2$ automorphism that contains only six $2$-cycles, which we saw is not positively realizable.∎
@@
Below we determine the positively realizable subgroups of $\mathrm{Aut}(P(10,3))$. By Theorem 2.1 we have Proposition 8.4.
@@
$\nu_{1}$ $=(u_{1}u_{5})(u_{2}v_{5}v_{1}u_{4})(u_{3}v_{2}v_{8}v_{4})(u_{6}u_{10})(u_{7}v_{10}v_{6}u_{9})(u_{8}v_{7}v_{3}v_{9})$
$\nu_{2}$ $=(u_{1}u_{6})(u_{2}u_{7})(u_{3}v_{7})(u_{4}v_{10})(u_{5}u_{10})(u_{8}v_{2})(u_{9}v_{5})(v_{1}v_{6})(v_{3}v_{4})(v_{8}v_{9}).$
Let $\Gamma_{D4}$ be the embedding of $P(10,3)$ in a standardly embedded solid torus $T$ in $S^{3}$, illustrated in Figure 9, where the edges which are on $\partial T$ are green on the flat torus and the edges of $\Gamma_{D4}$ which go into the interior of $T$ are red. The edges $\overline{u_{10}u_{1}}$ and $\overline{u_{5}u_{6}}$ are diametrically opposed arcs on the core of the solid torus illustrated in the center. The vertices in a given orbit of $\nu_{1}$ are the same color. Two special meridians are colored purple and orange on the right.
Figure 9. On the left and center, is an embedding $\Gamma_{D_{4}}$ of $P(10,3)$ on the boundary and core of a standardly embedded solid torus such that $\mathrm{TSG}_{+}(\Gamma_{D_{4}})=D_{4}$. On the right, we display two special meridians.
@@
Proposition 8.6.
Proof.
Figure 10. The $4$-simplex on the left is $\Omega$. The red graph in the center is the embedding $\Lambda$ of $P(10,3)$ in an unfolded picture of $\Omega$ such that $\mathrm{TSG}_{+}(\Lambda)=A_{5}\times\mathbb{Z}_{2}$. On the right is the $4$-simplex $\Omega^{\prime}$ which is dual to $\Omega$.
Since $h$ commutes with every element of the group $G$ of isometries, $H=\langle G,h\rangle=A_{5}\times\mathbb{Z}_{2}$, and $H$ leaves $\Lambda$ setwise invariant. Hence $A_{5}\times\mathbb{Z}_{2}\leq\mathrm{TSG}_{+}(\Lambda)$. Now since $A_{5}\times\mathbb{Z}_{2}$ is a maximal subgroup of $S_{5}\times\mathbb{Z}_{2}$ and $S_{5}\times\mathbb{Z}_{2}$ is not positively realizable for $P(10,3)$, we have $A_{5}\times\mathbb{Z}_{2}=\mathrm{TSG}_{+}(\Lambda)$. Thus $A_{5}\times\mathbb{Z}_{2}$ is positively realizable.
@@
Figure 11. The embedding $\Gamma_{D6}$ of $P(10,3)$ such that $\mathrm{TSG}_{+}(\Gamma_{D6}^{\prime})=D_{6}$, where $v_{10}$ is the point at $\infty$.
@@
Recall from the proof of Proposition 8.8 that $\Lambda_{T}$ was obtained from $\Lambda$ by adding a $4_{1}$ knot to the edges going between the tetrahedron $\Omega_{T}$ with vertices $1,2,3,4$ and the dual tetrahedron $\Omega^{\prime}_{T}$ with vertices $1245,1345,1235,1245$. Also $\mathrm{TSG}_{+}(\Lambda_{T})$ is induced by the group $A_{4}\times\mathbb{Z}_{2}$ of orientation preserving isometries of $(S^{3},\Omega_{T}\cup\Omega^{\prime}_{T})$. Now there is a reflection of $(S^{3},\Omega_{T}\cup\Omega^{\prime}_{T})$ which pointwise fixes the sphere containing the vertices $u_{9},v_{10},u_{1},u_{5}$ and $u_{8},v_{7},u_{2},u_{6}$. Thus the full group of isometries of $(S^{3},\Omega_{T}\cup\Omega^{\prime}_{T})$ including reflections is $S_{4}\times\mathbb{Z}_{2}$. Since these isometries induce a faithful action on $\Lambda_{T}$, it follows that $S_{4}\times\mathbb{Z}_{2}\leq\mathrm{TSG}(\Lambda_{T})$.
Let $\alpha=(u_{1}u_{2}u_{3}u_{4}u_{5}u_{6}u_{7}u_{8}u_{9}u_{10})(v_{1}v_{2}v_{3}v_{4}v_{5}v_{6}v_{7}v_{8}v_{9}v_{10})$. Then $\alpha\in\mathrm{Aut}(P(10,3))$ taking the edge $\overline{v_{1}v_{8}}$ to the edge $\overline{v_{2}v_{9}}$. However, in the embedding $\Lambda_{T}$, the edge $\overline{v_{1}v_{8}}$ contains a knot while the edge $\overline{v_{2}v_{9}}$ does not. Thus $\alpha$ is not contained in $\mathrm{TSG}(\Lambda_{T})$, and hence $\mathrm{TSG}(\Lambda_{T})\not=\mathrm{Aut}(P(10,3))=S_{5}\times\mathbb{Z}_{2}$. Since, $S_{4}\times\mathbb{Z}_{2}$ is a maximal subgroup of $S_{5}\times\mathbb{Z}_{2}$ it follows that $\mathrm{TSG}(\Lambda_{T})=S_{4}\times\mathbb{Z}_{2}$.
@@
We again start with the embedding $\Lambda_{T}$, but now we replace the $4_{1}$ knots on the edges $\overline{u_{7}v_{7}}$ and $\overline{u_{2}v_{2}}$ by the achiral and invertible knot $5_{2}$ to obtain an embedding $\Gamma_{D4}$ such that $\mathrm{TSG}(\Gamma_{D4})$ leaves $\{\overline{v_{7}u_{7}},\overline{u_{2}v_{2}}\}$ and $\{\overline{v_{8}v_{1}},\overline{u_{9}u_{10}},\overline{v_{6}v_{3}},\overline{u_{5}u_{4}}\}$ setwise invariant. Then any homeomorphism of $(S^{3},\Omega_{T}\cup\Omega_{T}^{\prime})$ which leaves $\Gamma_{D4}$ setwise invariant either interchanges the squares $S=\overline{1234}\subseteq\Omega_{T}$ and
@@
$\gamma_{1}$ $=(u_{5}v_{8}u_{9}v_{6})(u_{1}v_{10}u_{3}v_{4})(u_{7}v_{2})(v_{5}u_{8}v_{9}u_{6})(v_{1}u_{10}v_{3}u_{4})(u_{2}v_{7})$
$\gamma_{2}$ $=(u_{1}v_{10})(u_{2}v_{7})(u_{3}v_{4})(u_{6}v_{5})(u_{7}v_{2})(u_{8}v_{9})(v_{1}v_{3})(v_{6}v_{8})$
$\gamma_{3}$ $=(v_{3}v_{8})(v_{1}v_{6})(v_{9}v_{4})(v_{7}v_{2})(v_{5}v_{10})(u_{3}u_{8})(u_{1}u_{6})(u_{9}u_{4})(u_{7}u_{2})(u_{5}u_{10})$